Hardy type inequalities for Aharonov-Bohm magnetic potentials with multiple singularities

Alexander Balinsky School of Mathematics, Cardiff University, UK

The aim of this talk is to extend the Laptev-Weidl inequalities to the case of Aharonov-Bohm magnetic potentials with multiple singularities.